# Partition functions of determinantal point processes on polarized Kähler manifolds

### v3

@@ Abstract.
In this paper, we study the asymptotic expansion of the partition functions of determinantal point processes defined on a polarized Kähler manifold. The full asymptotic expansion of the partition functions is derived in two ways: one using Bergman kernel asymptotics and the other using the Quillen anomaly formula along with the asymptotic expansion of the Ray-Singer analytic torsion. By combining these two expressions, we show that each coefficient is given by geometric functionals on Kähler metrics satisfying the cocycle identity, and its first variation is closely related to the asymptotic expansion of the Bergman kernel. In particular, these functionals naturally generalize the Mabuchi functional in Kähler geometry and the Liouville functional on Riemann surfaces. Furthermore, we show that a Futaki-type holomorphic invariant obstructs the existence of critical points for each geometric functional given by the coefficients of the asymptotic expansion. Finally, we verify some of our results through explicit computations, which hold without the polarization assumption.
@@ 1. Introduction
In this paper, we study the asymptotic expansion of the partition functions of determinantal point processes defined on a polarized Kähler manifold. These processes were introduced by Berman in a series of papers [Ber13, Ber14, Ber17] and have been shown to satisfy a large deviation principle.
We derive a full asymptotic expansion of the partition function in two ways: one using Bergman kernel asymptotics and the other using the Quillen anomaly formula along with the asymptotic expansion of the Ray-Singer analytic torsion. By combining these two expressions, we analyze the properties of each coefficient in the asymptotic expansion. Each coefficient is given by geometric functionals on Kähler metrics that satisfy the cocycle identity, and its first variation can be expressed in terms of the coefficients of the Bergman kernel asymptotics. In particular, the second coefficient corresponds to the Mabuchi functional in Kähler geometry, while the third coefficient naturally generalizes the Liouville functional on Riemann surfaces to higher dimensions.
Furthermore, we show that the holomorphic invariant introduced by Futaki [Fut04] obstructs the existence of critical points for each geometric functional given by the coefficients of the asymptotic expansion. Although the construction relies on the polarization of the Kähler class, we demonstrate through direct computation that the results remain valid up to the third coefficient, without the polarization assumption. We will also discuss the connection between our results and the physics literature.
@@ 1.1. Mathematical background
Let $(M,\omega)$ be a polarized Kähler manifold. That is, there is a Hermitian line bundle $(L,h)$ over $M$ such that its Chern curvature form is given by $-i\omega$. Let $d_{k}:=\dim H^{0}(M,L^{k})$. The determinantal point process (DPP) associated with $L^{k}$ is a probability measure $\mu_{k,\omega}$ on $M^{d_{k}}$ defined as
$\mu_{k,\omega}:=\frac{1}{Z_{k}[h]}|\Psi^{k}(z_{1},\cdots,z_{d_{k}})|^{2}_{(h^{k})^{\boxtimes d_{k}}}\prod_{i=1}^{d_{k}}\frac{\omega^{n}}{n!},$ (1.1)
@@
where $\Psi^{k}$ is a holomorphic section of $(L^{k})^{\boxtimes d_{k}}$ over $M^{d_{k}}$ given by the Slater determinant. A more detailed definition will be given in Section 3. It was first introduced by Berman and subsequently studied by him, particularly regarding its relation to canonical Kähler metrics [Ber18, Ber22, Ber24]. See also [Fuj16, FO18, Aoi24] for related results.
@@
The first term in the large $k$ asymptotics of the normalization factor $Z_{k}[h]$, also known as the partition function, was obtained and proved to be equal to the Aubin-Yau functional by Berman-Boucksom [BB10]. In fact, they considered a more general setting, and their results have been recently extended in [Fin25]. In this paper, we focus on the (smooth) polarized case and derive the full asymptotic expansion of $Z_{k}[h]$ as $k\rightarrow\infty$, identifying its coefficients with geometric functionals on Kähler metrics.
@@
Note that $Z_{k}[h]$ is essentially the $\mathcal{L}$-functional introduced in [Don05], where its relation to the Bergman kernel asymptotics was also discussed. Following [KMMW17, SY25], we have an alternative way to obtain an asymptotic expansion of $Z_{k}[h]$ using the Quillen anomaly formula [BGS88] and the asymptotic expansion of the Ray-Singer analytic torsion [BV89, Fin18]. We will review the Quillen anomaly formula and Ray-Singer analytic torsion in Section 2.3. Combining these two approaches, we can study the properties of the coefficients in the asymptotic expansion.
@@
From the Quillen anomaly formula approach, the coefficients are expressed as geometric functionals involving Bott-Chern forms. Bott-Chern forms were first used by Donaldson [Don85, Don87] to construct geometric functionals and were further developed by Tian [Tia94, Tia00a] in the context of Kähler geometry. See also [Wei02]. Bott-Chern forms and related materials will be reviewed in Section 2.2.
@@
From the Bergman kernel asymptotics approach, the first variation of the geometric functionals is expressed in terms of integrals of scalar quantities involving the curvature of the Kähler metrics, arising from the Bergman kernel asymptotic expansion. In particular, the critical point equation is of the form $a_{j}-\Delta a_{j-1}\equiv\text{constant}$, where $a_{j}$ is the $j^{th}$ coefficient in the Bergman kernel asymptotics; see (2.2). Since the works of Catlin, Ruan, Tian, and Zelditch [Cat99, Rua98, Tia90, Zel98], Bergman kernel asymptotics has become a well-studied subject with various approaches; see, for example, [DLM06, BBS08, DK10, Xu12, HKSX16]. Especially, we will use the explicit expression for $a_{j},\;j\leq 3$ that Lu computed in [Lu00]. By the result of [LT04], we know that the equation $a_{j}-\Delta a_{j-1}\equiv\text{constant}$ is an elliptic equation of order $22j+$ 2 in Kähler potential. In fact, our geometric functionals $S_{j}$ (see Theorem 3.2 for the definition) can be regarded as the functionals they sought to find in [LT04].
@@
In [Fut04], Futaki introduced a family of holomorphic invariants that generalize various integral invariants, including the Futaki invariant [Fut83] and the Bando-Futaki invariant [Ban06]. We show that these invariants serve as obstructions to the existence of critical points of our geometric functionals $S_{j}$. As a byproduct, we extend Lu’s formula $(4.4)$ in [Lu04] to all $j\geq 0$, where it was originally verified by direct computation up to $j=2$.
@@
Throughout this paper, we emphasize when the polarization assumption is not used. In particular, in the Appendix A, we directly verify some of our results without the polarization assumption, following the spirit of original work of Mabuchi [Mab86].
@@ 1.2. Relation to physics literature
In the physics literature, the determinantal point process on Riemann surfaces corresponds to the plasma analogy in the quantum Hall effect (QHE). For a physics background on QHE, see [Ton16]. The form of asymptotic expansion of the corresponding partition function was conjectured in [ZW06] and later proved in [Kle14, KMMW17, SY25] for a pure bulk situation. For mathematical results in the presence of an edge, see [Ser24, BKS23, BKSY25] and references therein.
Recently, higher-dimensional QHE has also been studied in cases where the even-dimensional spatial manifold has a complex structure [KN16, KN23, AKN25]. In particular, an effective action for higher-dimensional QHE was derived in terms of Chern-Simons forms. We provide an alternative derivation of this result in Section 6. Note that in dimension $n=1$, [KMMW17] derived the effective action in terms of the Chern-Simons functional considering a more general moduli space.
In [Kle14], based on direct computations in dimension $n=1$, Klevtsov conjectured the form of the asymptotic expansion of partition function in all dimensions. We confirm his conjecture in Section 3 (see Remark 3). We also briefly discuss the relationship between the asymptotics of the QHE partition function and the 2D quantum gravity model in Section 5. Since this discussion is independent of the other parts of the paper, we defer it to Section 5.
@@ 1.3. Statement of the main results
Now we state our main results, starting with the asymptotic expansion of the DPP partition function. Note that in the main text, we will use a slightly different normalization convention, which will introduce additional $2\pi$ factors. Let $\mathcal{K}_{0}$ be the space of Kähler forms in $[\omega]$ and $\mathcal{K}_{\omega}$ be the space of Kähler potentials.
@@
    $S_{j}[\omega_{1},\omega_{0}]$ $=-S_{j}[\omega_{0},\omega_{1}],$ (1.3)
    $S_{j}[\omega_{2},\omega_{0}]$ $=S_{j}[\omega_{2},\omega_{1}]+S_{j}[\omega_{1},\omega_{0}];$ (1.4)
@@
    $\widehat{a_{j}(\omega_{\varphi})}:=\frac{1}{V}\int_{M}a_{j}(\omega_{\varphi})\frac{\omega_{\varphi}^{n}}{n!}=\frac{1}{V}\int_{M}\operatorname{Td}_{j}(T^{1,0}M)\operatorname{ch}_{n-j}(L),$ (1.6)
@@
    $S_{j}[\omega_{2},\omega_{1}]=s_{j}[\omega_{2}]-s_{j}[\omega_{1}],$ (1.7)
@@
$+\frac{i}{V}\left(\int_{M}\operatorname{Td}_{2}(T^{1,0}M)\frac{\omega^{n-2}}{(n-2)!}\right)\times\int_{M}\frac{-i}{(n+1)!}\sum^{n}_{s=0}\varphi\omega_{\varphi}^{s}\wedge\omega^{n-s}.$ (1.8)
As we will see, $S_{2}$ naturally generalizes the Liouville action to all dimensions. We refer to $S_{2}$ as the generalized Liouville action and compute its second variation as follows:
@@
$\;-\frac{1}{2}\Delta\dot{\varphi}\langle i\partial\bar{\partial}\dot{\varphi},\operatorname{Ric}\rangle\frac{\omega^{n}}{n!}+\frac{1}{12}\left(g^{i\bar{q}}g^{p\bar{j}}g^{r\bar{l}}g^{k\bar{s}}R_{p\bar{q}r\bar{s}}\partial_{i}\partial_{\bar{j}}\dot{\varphi}\partial_{k}\partial_{\bar{l}}\dot{\varphi}\right)\frac{\omega^{n}}{n!}.$ (1.9)
The following Theorem generalizes formula [Lu04, (4.4)] for all $j\geq 0$. Let $X$ be a holomorphic vector field on $M$ and $\theta_{X}$ be a purely imaginary function satisfying $\iota_{X}\omega=-\bar{\partial}\theta_{X}$.
@@
Let $(M,\omega)$ be a polarized Kähler manifold. Let $X$ and $\theta_{X}$ be given as above. Then for all $j\geq 0$, the following identity holds :
@@
Under the normalization of $\theta_{X}$ by $\int_{M}\theta_{X}\omega^{n}=0$, it is independent of the choice of the Kähler metric in $\operatorname{c}_{1}(L)$.
The RHS of (1.10) is precisely the holomorphic invariant introduced in [Fut04]. Note that these invariants obstruct the existence of critical points of $S_{j}$. When $\mathfrak{h}(M)=0$, these obstructions vanish. In this context, assuming that $Aut(M,L)$ is discrete, we can prove the following proposition using Donaldson’s result [Don01] on balanced metrics.
@@
Finally, we derive a formula for the effective action of the higher-dimensional QHE, as presented in [KN16], expressed in terms of Chern-Simons forms.
@@
$2\int_{M\times[0,1]}CS(\operatorname{ch}_{n+1};\bm{\nabla}_{\mathbf{L}},\bm{\nabla^{0}}_{\mathbf{L}}).$ (1.12)
@@ 1.4. Structure of the paper
Acknowledgements: The author would like to thank Siarhei Finski for clarifications regarding his Theorem and Remark 1. This research was supported by NRF Grant RS-2024-00346651.

### v5

@@ Abstract.
In this paper, we study the full asymptotic expansion of the partition functions of determinantal point processes defined on a polarized Kähler manifold. We show that the coefficients of the expansion are given by geometric functionals on Kähler metrics satisfying the cocycle identity, whose first variations can be expressed through the TYZ expansion coefficients of the Bergman kernel. In particular, these functionals naturally generalize the Mabuchi functional in Kähler geometry and the Liouville functional on Riemann surfaces. We further show that Futaki-type holomorphic invariants obstruct the existence of critical points of these geometric functionals, extending Lu’s formula. We also verify that certain formulas remain valid up to the third coefficient without assuming polarization. Finally, we discuss the relation of our results to the quantum Hall effect (QHE), where the determinantal point process provides a microscopic model. In particular, we recover the higher-dimensional effective Chern-Simons actions derived in the physics literature and confirm a conjecture of Klevtsov on the form of the partition function asymptotics.
@@ 1. Introduction
In this paper, we study the asymptotic expansion of the partition functions of determinantal point processes defined on a polarized Kähler manifold. These processes were introduced by Berman in a series of papers [berman2013determinantal, berman2014determinantal, berman2017large] and have been shown to satisfy a large deviation principle.
We derive a full asymptotic expansion of the partition function in two ways: one using Bergman kernel asymptotics and the other using the Quillen anomaly formula along with the asymptotic expansion of the Ray-Singer analytic torsion. By combining these two expressions, we analyze the properties of each coefficient in the asymptotic expansion. Each coefficient is given by geometric functionals on Kähler metrics that satisfy the cocycle identity, and its first variation can be expressed in terms of the coefficients of the Bergman kernel asymptotics. In particular, the second coefficient corresponds to the Mabuchi functional in Kähler geometry, while the third coefficient generalizes the Liouville functional on Riemann surfaces to higher dimensions.
Furthermore, we show that the holomorphic invariant introduced by Futaki [futaki2004asymptotic] obstructs the existence of critical points for each geometric functional given by the coefficients of the asymptotic expansion. Although the construction relies on the polarization of the Kähler class, we demonstrate through direct computation that the results remain valid up to the third coefficient, without the polarization assumption. We will also discuss the connection between our results and the physics literature.
@@ 1.1. Mathematical background
Let $(M,\omega)$ be a polarized Kähler manifold. That is, there is a Hermitian line bundle $(L,h)$ over $M$ such that its Chern curvature form is given by $i\omega$. Let $d_{k}:=\dim H^{0}(M,L^{k})$. The determinantal point process (DPP) associated with $L^{k}$ is a probability measure $\mu_{k,\omega}$ on $M^{d_{k}}$ defined as
$\mu_{k,\omega}:=\frac{1}{Z_{k}[h]}|\Psi^{k}(z_{1},\cdots,z_{d_{k}})|^{2}_{(h^{k})^{\boxtimes d_{k}}}\prod_{i=1}^{d_{k}}\frac{\omega^{n}}{n!},$
@@
where $\Psi^{k}$ is a holomorphic section of $(L^{k})^{\boxtimes d_{k}}$ over $M^{d_{k}}$ given by the Slater determinant. A more detailed definition will be given in Section 3. It was introduced by Berman and subsequently studied by him, particularly regarding its relation to canonical Kähler metrics [berman2018kahler, berman2022measure, berman2024probabilistic]. See also [fujita2016berman, fujita2018k, aoi2024microscopic] for related results.
@@
The normalization factor $Z_{k}[h]$, also known as the partition function, is essentially Donaldson’s $\mathcal{L}_{k}$-functional introduced in [donaldson2005scalar]. The first term in the large $k$ asymptotics of $Z_{k}[h]$ without assuming positivity of the curvature of $h$, was studied by Berman-Boucksom [berman2010growth], extending the method of [donaldson2005scalar]. For a recent extension of their results, see [finski2025small]. In this paper, we focus on the (smooth) positive curvature case and derive the full asymptotic expansion of $Z_{k}[h]$ as $k\rightarrow\infty$, identifying its coefficients with geometric functionals on Kähler metrics.
@@
The relation between the asymptotic expansion of $Z_{k}[h]$ and Bergman kernel asymptotics was given in [donaldson2005scalar]. Following [klevtsov2017quantum, shen2025geometric], we have an alternative way to obtain an asymptotic expansion of $Z_{k}[h]$ using the Quillen anomaly formula [bismut1988analytic] and the asymptotic expansion of the Ray-Singer analytic torsion [vasserot1989asymptotics, finski2018full]. We will review the Quillen anomaly formula and Ray-Singer analytic torsion in Section 2.3. Combining these two approaches, we can study the properties of the coefficients in the asymptotic expansion.
@@
From the Quillen anomaly formula approach, the coefficients are expressed as geometric functionals involving Bott-Chern forms. Bott-Chern forms were first used by Donaldson [donaldson1985anti, donaldson1987infinite] to construct geometric functionals and were further developed by Tian [tian1994k, tian1999bott] in the context of Kähler geometry. See also [weinkove2002higher]. Bott-Chern forms and related materials will be reviewed in Section 2.2.
@@
From the Bergman kernel asymptotics approach, the first variation of the geometric functionals is expressed in terms of integrals of scalar quantities involving the curvature of the Kähler metrics, arising from the Bergman kernel asymptotic expansion. In particular, the critical point equation is of the form $a_{j}-\Delta a_{j-1}\equiv\text{constant}$, where $a_{j}$ is the $j^{th}$ coefficient in the TYZ expansion of the Bergman kernel; see (2.1). Since the works of Catlin, Ruan, Tian, and Zelditch [catlin1999bergman, ruan1996canonical, tian1990set, zelditch1998szego], Bergman kernel asymptotics has become a well-studied subject with various approaches; see, for example, [dai2006asymptotic, berman2008direct, douglas2010bergman, xu2012closed, hezari2016asymptotic]. Especially, we will use the explicit expression for $a_{j},\;j\leq 3$ that Lu computed in [lu2000lower]. By the result of [lu2004log], we know that the equation $a_{j}-\Delta a_{j-1}\equiv\text{constant}$ is an elliptic equation of order $2j+2$ in Kähler potential.
@@
In [futaki2004asymptotic], Futaki introduced a family of holomorphic invariants that generalize various integral invariants, including the Futaki invariant [futaki1983compact] and the Bando-Futaki invariant [bando2006obstruction]. We show that these invariants serve as obstructions to the existence of critical points of our geometric functionals $S_{j}$. As a byproduct, we extend Lu’s formula $(4.4)$ in [lu2004k] to all $j\geq 0$, where it was originally verified by direct computation up to $j=2$.
In the standard GIT picture of Kähler geometry, geometric functionals correspond to the log-norm functionals, while Futaki-type invariants correspond to the weight. See also [foth2007manifold, eum2025asymptotic] for a moment map interpretation of the functions $a_{j}-\Delta a_{j-1}$. Thus, equations of the form $a_{j}-\Delta a_{j-1}\equiv\text{constant}$ formally fit into the GIT setting except the convexity of the log-norm functional. We compute the second variation of $S_{2}$ with a view to this problem. See [cristofori2025third] and references therein for related work on equations of the form $a_{j}\equiv\text{constant}$.
@@
Throughout this paper, we emphasize when the polarization assumption is not used. In particular, in the Appendix A, we directly verify some of our results without the polarization assumption, following the spirit of original work of Mabuchi [mabuchi1986k].
@@ 1.2. Relation to physics literature
In the physics literature, the determinantal point process on Riemann surfaces corresponds to the plasma analogy in the quantum Hall effect (QHE). For a general physics background on the QHE, see [tong2016lectures]; for the role of geometric functionals in the QHE, see [can2015geometry, klevtsov2016geometry]. The form of asymptotic expansion of the corresponding partition function was conjectured by Zabrodin-Wiegmann in [zabrodin2006large] for $\mathbb{CP}^{1}$ and later proved in [klevtsov2014random, klevtsov2017quantum, shen2025geometric] for arbitrary Riemann surfaces with pure bulk assumption. For mathematical results in the presence of an edge, see [serfaty2024lectures, byun2023partition, byun2025free] and references therein.
Recently, higher-dimensional QHE has also been studied in cases where the even-dimensional spatial manifold has a complex structure [karabali2016geometry, karabali2023transport, agarwal2025fractional]. In particular, an effective action for higher-dimensional QHE was derived in terms of Chern-Simons forms. We provide an alternative derivation of this result in Section 6. In dimension $n=1$, Klevtsov-Ma-Marinescu-Wiegmann [klevtsov2017quantum] derived the effective action in terms of the Chern-Simons functional considering a more general moduli space.
In [klevtsov2014random], based on direct computations in dimension $n=1$, Klevtsov conjectured the form of the asymptotic expansion of partition function in all dimensions. We confirm his conjecture in Section 3 (see Remark 3). We also briefly discuss the relationship between the asymptotics of the QHE partition function and the 2D quantum gravity model in Section 5. Since this discussion is independent of the other parts of the paper, we defer it to Section 5.
@@ 1.3. Statement of the main results
Now we state our main results, starting with the asymptotic expansion of the DPP partition function. Note that in the main text, we will use a slightly different normalization convention, which will introduce additional $2\pi$ factors. Also we will identify a hermitian metric $h$ with Kähler potential $\varphi$. Let $\mathcal{K}_{0}$ be the space of Kähler forms in $\omega]$ and $\mathcal{K}_{\omega}$ be the space of Kähler potentials.
@@
    $S_{j}[\omega_{1},\omega_{0}]$ $=-S_{j}[\omega_{0},\omega_{1}],$
    $S_{j}[\omega_{2},\omega_{0}]$ $=S_{j}[\omega_{2},\omega_{1}]+S_{j}[\omega_{1},\omega_{0}];$
@@
    $\widehat{a_{j}(\omega_{\varphi})}:=\frac{1}{V}\int_{M}a_{j}(\omega_{\varphi})\frac{\omega_{\varphi}^{n}}{n!}=\frac{1}{V}\int_{M}\operatorname{Td}_{j}(T^{1,0}M)\operatorname{ch}_{n-j}(L),$
@@
    $S_{j}[\omega_{2},\omega_{1}]=s_{j}[\omega_{2}]-s_{j}[\omega_{1}],$
@@
$+\frac{i}{V}\left(\int_{M}\operatorname{Td}_{2}(T^{1,0}M)\frac{\omega^{n-2}}{(n-2)!}\right)\times\int_{M}\frac{-i}{(n+1)!}\sum^{n}_{s=0}\varphi\omega_{\varphi}^{s}\wedge\omega^{n-s}.$
As we will see, $S_{2}$ reduces to the Liouville action in dimension $n=1$. We compute its second variation as follows:
@@
$\;-\frac{1}{2}\Delta\dot{\varphi}\langle i\partial\bar{\partial}\dot{\varphi},\operatorname{Ric}\rangle\frac{\omega^{n}}{n!}+\frac{1}{12}\left(g^{i\bar{q}}g^{p\bar{j}}g^{r\bar{l}}g^{k\bar{s}}R_{p\bar{q}r\bar{s}}\partial_{i}\partial_{\bar{j}}\dot{\varphi}\partial_{k}\partial_{\bar{l}}\dot{\varphi}\right)\frac{\omega^{n}}{n!}.$
The following Theorem generalizes formula [lu2004k, (4.4)] for all $j\geq 0$. Let $X$ be a holomorphic vector field on $M$ and $\theta_{X}$ be a holomorphy potential function satisfying $\iota_{X}\omega=-\bar{\partial}\theta_{X}$.
@@
Let $(M,\omega)$ be a polarized Kähler manifold. Let $X$ and $\theta_{X}$ be given as above and suppose it is purely imaginary. Then for all $j\geq 0$, we have the following identity:
@@
More generally, for non purely imaginary $\theta_{X}$, we have
$\int_{M}i{\rm Im\,}\theta\left(a_{j}(\omega)-\Delta a_{j-1}(\omega)\right)\frac{\omega^{n}}{n!}=\frac{1}{(n+1-j)!}\int_{M}\operatorname{Td}_{j}(R+\frac{1}{2}(\nabla X+(\overline{\nabla X})^{\flat,\sharp}))(\omega+i{\rm Im\,}\theta)^{n+1-j}.$ (1.4)
Under the normalization of $\theta_{X}$ by $\int_{M}\theta_{X}\omega^{n}=0$, the right hand sides of (1.3) and (1.4) are independent of the choice of the Kähler metric in $\operatorname{c}_{1}(L)$.
The right hand side of (1.3) is precisely the holomorphic invariant introduced in [futaki2004asymptotic]. We will show that these invariants obstruct the existence of critical points of $S_{j}$. When there is no holomorphic vector fields, these obstructions vanish. In this context, assuming that $Aut(M,L)$ is discrete, we can prove the following proposition using Donaldson’s result [donaldson2001scalar] on balanced metrics.
@@
Finally, we derive a formula for the effective action of the higher-dimensional QHE, as presented in [karabali2016geometry], expressed in terms of Chern-Simons forms.
@@
$2\int_{M\times[0,1]}CS(\operatorname{ch}_{n+1};\bm{\nabla}_{\mathbf{L}},\bm{\nabla^{0}}_{\mathbf{L}}).$
@@ 1.4. Structure of the paper
Acknowledgements: The author would like to thank Siarhei Finski for clarifications regarding his Theorem and Remark 1. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) RS-2024-00346651.